# Decision-Making Behavior Evaluation Framework for LLMs under Uncertain Context

### v1

@@ 6 Discussion
Implications for developers and researchers: Previous research [16] has demonstrated that biases can be injected at various levels using different instructions, and our experiments provide evidence that injected demographic features can significantly affect how LLMs make decisions. Critical questions arise: should LLMs mirror human decision-making processes, including preexisting biases, or aim to satisfy ethical standards that remediate these flaws? How should LLMs perform when assisting in human decision-making? Prejudices objectively exist in human society, which are inevitably absorbed by the LLM from their datasets, training, and even aligning with Reinforcement Learning from Human Feedback (RLHF), but LLMs may still raise the bar for ethics by locating potential biases and addressing them carefully. Answering these questions involve balancing the usability and relatability of LLMs, with the ethical responsibility to avoid perpetuating harmful biases. While mimicking human behavior can make LLMs more realistic and relatable, correcting biases can lead to fairer, more equitable decision-making processes and avoid reinforcing existing prejudices.
Our study presents intriguing discrepancies between human and LLM behaviors. For example, while human studies may not find significant differences in risk preferences among sexually minority groups [24], our LLM results suggest otherwise. This discrepancy may result from two potential reasons: 1. Flaws in Human Studies: Studies on minority groups often face privacy challenges in obtaining representative samples for their personal information. 2. LLM misunderstanding: LLMs may possess flawed understandings due to biases from data and training. If LLMs could reflect real-world trends more accurately from broader and cleaner data and more careful training, should researchers consider LLM outputs as supplementary evidence for social studies? It can only be answered when further studies can address these discrepancies between human and LLM data.
@@
Limitations: One limitation of this work is the difficulty in directly comparing LLM behaviors to humans due to the complex and sensitive nature of many demographic features, especially those related to minority groups, making it challenging for human subjects to process in studies. Additionally, while this study uses financial experiments to elicit overall preferences, exploring other domains with other experiments could provide a more comprehensive understanding of LLM behaviors.
@@ 7 Conclusion
In conclusion, our work establishes a foundational framework for evaluating LLM behaviors and opens avenues for future research aimed at aligning these models more closely with ethical standards and human values. We evaluated three LLMs regarding their capability to process decision-making decisions through three perspectives: risk, probability, and loss. While all models exhibit general tendencies akin to human behavior, each demonstrates unique deviations to an extent. The incorporation of socio-demographic features into our analysis revealed significant impacts of human-related features on the LLM decision-making process, underscoring the potential biases and variability in their outputs. By addressing the complexities and biases inherent in LLM decision-making, we can better harness their capabilities to support fair and equitable outcomes in real-world applications. These findings emphasize the critical need for ongoing scrutiny and refinement of LLMs to ensure they do not perpetuate or exacerbate societal biases. Additionally, our work raises further exploration about how LLMs should be designed to balance realism with ethical responsibility, and what their intrinsic behaviors should reflect, with and without human-demographic embeddings. This marks the boundary between closed-loop answering systems and genuine artificial intelligence.

### v2

@@ 6 Discussion
Implications for developers and researchers: Previous research [15] has demonstrated that biases can be injected at various levels using different instructions, and our experiments provide evidence that injected demographic features can significantly affect how LLMs make decisions. Critical questions arise: should LLMs mirror human decision-making processes, including preexisting biases, or aim to satisfy ethical standards that remediate these flaws? How should LLMs perform when assisting in human decision-making? Prejudices objectively exist in human society, which are inevitably absorbed by the LLM from their datasets, training, and even aligning with Reinforcement Learning from Human Feedback (RLHF), with advancements such as Chain-of-Thought (CoT) prompting [35], LLMs could potentially raise the bar for ethical decision-making by helping identify and correct biases through more structured reasoning. Moreover, LLMs that leverage these structured reasoning frameworks could potentially distinguish between decisions that reflect societal biases and those that follow ethical standards, prompting users to reconsider biased responses. If biases are corrected, methods such as targeted fine-tuning and RLHF adjustments could be used to realign the model’s behavior toward fairness. However, a key question remains: in doing so, do we risk making LLMs less reflective of human-like behavior, and how do we balance ethical responsibility with the usability of these models in real-world applications?
Our study presents intriguing discrepancies between human and LLM behaviors. For example, while human studies may not find significant differences in risk preferences among sexually minority groups [3], our LLM results suggest otherwise. This discrepancy may result from two potential reasons: 1. Flaws in Human Studies: Studies on minority groups often face privacy challenges in obtaining representative samples for their personal information. 2. LLM misunderstanding: LLMs may possess flawed understandings due to biases from data and training. If LLMs could reflect real-world trends more accurately from broader and cleaner data and more careful training, should researchers consider LLM outputs as supplementary evidence for social studies? This can only be answered when further studies can address these discrepancies between human and LLM data.
@@
Limitations: One limitation of this work is the difficulty in directly comparing LLM behaviors to humans due to the complex and sensitive nature of many demographic features, especially those related to minority groups, making it challenging for human subjects to process in studies. Additionally, while this study uses financial experiments to elicit overall preferences, exploring other domains with other experiments could provide a more comprehensive understanding of LLM behaviors. We recommend future work explore multiple models to compare model fit and robustness for a deeper understanding of LLM decision-making.
@@ 7 Conclusion
In conclusion, our work establishes a foundational framework for evaluating LLM behaviors and opens avenues for future research aimed at aligning these models more closely with ethical standards and human values. We evaluated three LLMs regarding their capability to process decision-making decisions through three perspectives: risk, probability, and loss. While all models exhibit general tendencies akin to human behavior, each demonstrates unique deviations to an extent. The incorporation of socio-demographic features into our analysis revealed significant impacts of human-related features on the LLM decision-making process, underscoring the potential biases and variability in their outputs. By addressing the complexities and biases inherent in LLM decision-making, we can better harness their capabilities to support fair and equitable outcomes in real-world applications. These findings emphasize the critical need for ongoing scrutiny and refinement of LLMs to ensure they do not perpetuate or exacerbate societal biases. Additionally, our work raises further exploration about how LLMs should be designed to balance realism with ethical responsibility and what their intrinsic behaviors should reflect, with and without human-demographic embeddings. This marks the boundary between closed-loop answering systems and genuine artificial intelligence.